Determine the inverse Laplace transform of the function:

F(s) = [10(s + 3)(s + 5)]/[s(s + 1)(s + 6)]

and build a plot as a function of time.

Step by step neatly please. thanks
Answer #1

Given function: F(S): 1005+3)(5+5) s(s+) (+6) A + B + 1. = SH S+6 A = F(S) = locow 1000+3) (0+5) (ot) cotes 104345-25 lx6 1s=